OpenStudy (meilendurcer):

11 1/3 ÷ 4 1/4= write out

OpenStudy (anonymous):

first convert the given mixed numbers to improper fractions then multiply the first by the multiplicative inverse of second cancel out common factors and get the answer

OpenStudy (anonymous):

So first work out how many thirds in 11 1/3, then how many quarters in 4 1/4

OpenStudy (anonymous):

11 1/3 = 34/3 and 4 1/4 = 17/4

OpenStudy (meilendurcer):

thanks. !